OpenStudy (anonymous):

Find the exponential function f(x)=Ca^x whose graph goes through the points (0,4) and (2,16).

OpenStudy (across):

My beloved friend, I would help you, but I am a Barcelona fan. :P Sincerely, Across

OpenStudy (anonymous):

aww cmon

OpenStudy (anonymous):

although im a madrid fan, i respect barca

OpenStudy (anonymous):

lol

OpenStudy (across):

I will give you a hint: When x = 0, then we have\[f(x)=Ca^{0}=C.\]Therefore, C should be 4 such that we obtain the point (0, 4). What should a then be?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Try substituting into your expression the values for x and y and see what that tells you. For instance, if (0,4) is on the graph of f(x)=Ca^x Then f(0) = 4 i.e., 4 = Ca^0 = C Now try the other point.

OpenStudy (anonymous):

thanks for the help hala madrid respect barca lol

OpenStudy (anonymous):

also thank you simons

OpenStudy (across):

You're correct, Jinnie. ;)

myininaya (myininaya):

are these baseball teams?

OpenStudy (anonymous):

no, football

OpenStudy (anonymous):

phillies start in ten minutes!

myininaya (myininaya):

i don't do sports

OpenStudy (across):

You don't know what these teams are?! Jinnie, I think we should get together and linch myininaya. ;P

myininaya (myininaya):

lol

OpenStudy (anonymous):

\[f(0)=C\times a^0=C=4\] soy ou have \[f(x)=4\times a^x\] now since \[f(2)=4\times a^2=16\] you get \[a^2=4\implies a = 2\]

OpenStudy (anonymous):

*linch = lynch

OpenStudy (anonymous):

good job guys, you were all helpful i dont know who to give best answer to

myininaya (myininaya):

you can give to all three: satellite,across, and simmons

OpenStudy (anonymous):

and the phillies. give them a medal as well

myininaya (myininaya):

phillies will lose i'm against them just because satellite is for them

OpenStudy (anonymous):

either i am making typos galore, or site is messing up my text...

OpenStudy (anonymous):

has baseball made it to arkansas yet?

myininaya (myininaya):

lol

myininaya (myininaya):

i don't think arkansas has any professional teams

myininaya (myininaya):

i could be wrong

OpenStudy (anonymous):

a google search finds "arkansas travelers" a minor league team. in truth sports escapes me.